OpenStudy (anonymous):

The following functions f and g are periodic. You can get the graph of g by changing the amplitude and period of f, and then applying a translation. Find function g(x) in terms of f(x).

OpenStudy (anonymous):

Here are the graphs I have to work with.

OpenStudy (anonymous):

@myininaya, @quickstudent, @Compassionate, @Hero, @triciaal I need help on this

OpenStudy (anonymous):

@triciaal by any chance would there be a way of doing this because my issue is knowing where and what to start

OpenStudy (triciaal):

I am not really sure. what I am thinking of is finding the function f(x) and identifying the changes to get it to g(x). f(x) max (pi/2,1) min (3/2*pi, -1) ,(0, 0), (2*pi, 0), (5/2pi. 1)

OpenStudy (triciaal):

g(x) (3/4pi, 0), (7/2pi 0) max 3, min 0

OpenStudy (triciaal):

for a given point on f(x) look at the corresponding point on g and how it has "moved"

OpenStudy (triciaal):

minimum g(x)= 0

OpenStudy (anonymous):

although for the question would I have to show a new translated graph or just use the f(x) graph (the picture) and draw the transition?

OpenStudy (anonymous):

@campbell_st i need assistance

OpenStudy (campbell_st):

well f(x) = sin(x) g(x) has a period of pi... so the period is found using \[period = \frac{2\pi}{b}\] so b = 2 g(x) seems to have an amplitude of 3... and is centred on 1.5 so find mine g(x) = 3sin(2x) + 1.5 hope it helps

OpenStudy (aum):

I think the amplitude of g(x) is 1.5 and \(\large g(x) = 1.5f(2x) + 1.5\)

OpenStudy (campbell_st):

dang... amplitude should be 1.5 so I'd say g(x) = 1.5sin(2x) + 1.5

OpenStudy (anonymous):

so technically its just asking me to write an equation to find g(x).

OpenStudy (campbell_st):

well its asking you to find the equation of g(x) in terms of f(x)

OpenStudy (anonymous):

My first thought in my head was "I think I need to draw the translation for this."

OpenStudy (anonymous):

but basically this equation is only for g(x)

OpenStudy (campbell_st):

its a case of using your knowledge to get the equation... look at the period... and the amplitude... and the fact its been shifted up...

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@aaronq